OpenStudy (anonymous):

can someone tell me if this is prime, I can't figure it out x^3-729 for some reason I am stuck!!

OpenStudy (amistre64):

if 729 can be cube rooted, then this demonstrates a difference of cubes

OpenStudy (amistre64):

9*81 = 729 9*9*9 = 729; 9^3 = 729. this can be factored as a difference of cubes

OpenStudy (anonymous):

ok I get that part so now what do I need to do or how to I write it an answer form.

OpenStudy (amistre64):

i think the general set up is: x^3 - y^3 = (x-y) (x^2 +xy +y^2)

OpenStudy (amistre64):

but if the question is asking for prime or not; then its not prime since both terms are ... real cubics for lack of a better term

OpenStudy (anonymous):

well it says factor but if not then it is prime x^3-729=??

OpenStudy (amistre64):

then it factors into the general form; and keep an eye in the quadratic factor to see if its still factorable

OpenStudy (amistre64):

if y^2-4y^2 is negative then its prime across the reals